# Multimodal Self-Instruct: Synthetic Abstract Image and Visual Reasoning Instruction Using Language Model

### v4

@@ Abstract
Wenqi Zhang1,∗, Zhenglin Cheng1,∗, Yuanyu He1, Mengna Wang2, Yongliang Shen1 Zeqi Tan1, Guiyang Hou1, Mingqian He1, Yanna Ma3, Weiming Lu1, Yueting Zhuang1 1College of Computer Science and Technology, Zhejiang University 2Institute of Software, Chinese Academy of Sciences 3University of Shanghai for Science and Technology {zhangwenqi, luwm}@zju.edu.cn Project Page: https://multi-modal-self-instruct.github.io
11footnotetext: The first two authors have equal contributions.
@@ 1 Introduction
In recent times, spurred by breakthroughs in large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023a; OpenAI, 2022, 2023; Touvron et al., 2023b; Bi et al., 2024; Jiang et al., 2024; Anthropic, 2024; Abdin et al., 2024), large multimodal models (LMMs) have also undergone rapid advancements (Liu et al., 2024b, a; Team et al., 2023; Bai et al., 2023a; Lu et al., 2024; McKinzie et al., 2024). Leveraging a pre-trained LLM to encode all modalities empowers LMMs to understand human daily environments and execute complex tasks (Hong et al., 2023; Zhang et al., 2023b; Hu et al., 2023; Zhang et al., 2023a; Koh et al., 2024; Zhang et al., 2024c). This greatly expands the potential of general-purpose AI assistants.
@@ 3.3 Implementation Details
Data Selection Size Chart (%) Table (%) Map (%)
@@
Table 3: We investigate the synergistic effects between the three tasks. Chart and table corpus can improve each other and both benefit road map tasks.
@@ 4.1 Settings
We evaluated the performance of mainstream open-source and closed-source LMMs, including Llava-1.5-7B (Liu et al., 2024a), Llava-1.5-13B, InstructBLIP-7B (Dai et al., 2024), InstructBLIP-13B, Deepseek-VL-Chat-1.3B (Lu et al., 2024), Deepseek-VL-Chat-7B, Claude-3.5-Sonnet, Claude-3-Sonnet [^4] https://www.anthropic.com/news/claude-3-family, GPT-4o, GPT-4-Vision-1106 (OpenAI, 2023), Gemini-1.5-pro [^5] https://deepmind.google/technologies/gemini/pro/ and Qwen-VL-Plus (Bai et al., 2023b). All models were evaluated using the same prompts and temperature settings. We provide the evaluation metrics and other training details in  Appendix A.
@@ 4.4 Generalized to Untrained Tasks
4.5 How to improve the abstract image comprehension capabilities of LMMs?
More than just for instruction fine-tuning, our experiments have revealed that the abstract image comprehension capabilities of Large LMMs can be enhanced through various approaches:
Designing More Versatile Visual Encoders: First, we observe that current LMMs have weak visual representation abilities for abstract images, which may be caused by the current visual encoders. Most of them use clip-based encoders, which emphasize semantic features while neglecting purely visual features. We plan to explore adding another visual encoder, such as DinoV2, to Llava to improve its understanding of abstract images.
Increasing Image Resolution: Then we also observed that most current LMMs resize the original image to a resolution of 336x336 when training, as it reduces the number of visual tokens without losing much semantic information. However, for these abstract images composed of lines and geometric shapes, lowering the resolution results in the loss of a significant amount of geometric features, thereby affecting the ability of abstract visual perception. Increasing the image resolution and training LMMs from scratch may be a good solution.
Incorporate into Pre-training: Next, we will incorporate abstract image data during the pre-training stage, not just during the SFT stage. This is because, while benchmarking current LMMs, we found that their weakness lies in the perception of abstract images rather than in their instruction-following abilities.
@@ 4.5 How to improve the abstract image comprehension capabilities of LMMs?
We believe that focusing on the aforementioned four aspects can effectively enhance the abstract image comprehension capabilities of LMMs without compromising their general capabilities. We treat this as a direction for future work.
@@ 5 Conclusions
Our multi-modal strategy can synthesize a vast amount of abstract images and reasoning instructions, providing valuable training data to enhance LMMs. However, we want to highlight that there remain some limitations or areas for improvement: 1. Our data synthesis process relies on the code generation and reasoning capabilities of LLMs, which are only available in closed-source models like GPT-4. Using these models is costly. As the capabilities of open-source models improve, we are attempting to use open-source LLMs, such as Llama 3 and Deepseek-V2, to synthesize data. This will significantly reduce our expenses. 2. This work used code to synthesize abstract images in eight scenarios, such as tables and maps. In the future, we can expand to more scenarios, such as using code to control robot simulators to generate specific house layouts and structures, thereby producing a massive amount of data. 3. We believe that the image resolution of visual encoders is a bottleneck for current LMMs, especially for these abstract diagrams. In the future, we plan to improve the image resolution of the encoders to enhance the fine-grained perception capabilities of LMMs.
@@ D.3 Data Synthesis
Data synthesis is widely used in LLM training to supplement the insufficiency of instruction-following data. Many studies focus on generating high-quality synthetic data either distilling dialogue data from a strong LLM (Wang et al., 2022b; Xu et al., 2023a; Yu et al., 2023a; Chen et al., 2023a; Zhao et al., 2023), or using external tools to refine LLM-generated synthetic data (Wei et al., 2023; Lee et al., 2024). For instance, Wang et al. (2022b) proposed Self-Instruct to improve the instruction-following ability of LLMs via their own generation of instruction data. Xu et al. (2023a) further generated more complex instruction through Evol-Instruct. Yu et al. (2023a) synthesized a mathematical dataset from LLMs by bootstrapping mathematical questions and rewriting the question from multiple perspectives. Wei et al. (2023) can generate diverse and realistic coding problems from open-source code snippets. Lei et al. (2024) can also create high-quality large code datasets for LLMs. It simulates programmers writing code and conducting unit tests through agent interactions, ensuring annotation accuracy with an external code executor.

### v5

@@ Abstract
Wenqi Zhang1,∗, Zhenglin Cheng1,∗, Yuanyu He1, Mengna Wang2, Yongliang Shen1 Zeqi Tan1, Guiyang Hou1, Mingqian He1, Yanna Ma3, Weiming Lu1,†, Yueting Zhuang1 1College of Computer Science and Technology, Zhejiang University 2Institute of Software, Chinese Academy of Sciences 3University of Shanghai for Science and Technology {zhangwenqi, luwm}@zju.edu.cn Project Page: https://multi-modal-self-instruct.github.io
11footnotetext: The first two authors have equal contributions.22footnotetext: Corresponding author.
Figure 1: Benchmarking Leading LMMs on abstract image understanding and reasoning tasks.
@@ 1 Introduction
In recent times, spurred by breakthroughs in large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023a; OpenAI, 2022, 2023; Touvron et al., 2023b; Bi et al., 2024; Jiang et al., 2024; Anthropic, 2024; Abdin et al., 2024), large multimodal models (LMMs) have also undergone rapid advancements (Liu et al., 2024b, a; Team et al., 2023; Bai et al., 2023a; Lu et al., 2024; McKinzie et al., 2024). Leveraging a pre-trained LLM to encode all modalities empowers LMMs to understand human daily environments and execute complex tasks (Hong et al., 2023; Zhang et al., 2023b; Hu et al., 2023; Zhang et al., 2023a, 2024c; Koh et al., 2024; Zhang et al., 2024d). This greatly expands the potential of general-purpose AI assistants.
@@ 3.3 Implementation Details
Data Selection Size Chart Table Map
@@
Table 3: We investigate the synergistic effects between the three tasks (acc in %). Chart and table corpus can improve each other and both benefit road map tasks.
@@ 4.1 Settings
We evaluated the performance of mainstream open-source and closed-source LMMs, including Llava-1.5-7B (Liu et al., 2024a), Llava-1.5-13B, InstructBLIP-7B (Dai et al., 2024), InstructBLIP-13B, Deepseek-VL-Chat-1.3B (Lu et al., 2024), Deepseek-VL-Chat-7B, Claude-3.5-Sonnet, Claude-3-Sonnet, GPT-4o, GPT-4-Vision-1106 (OpenAI, 2023), Gemini-1.5-pro, MiniCPM-2.6 (Yao et al., 2024), and Qwen-VL-Plus (Bai et al., 2023b). All models were evaluated using the same prompts and temperature settings. We provide the evaluation metrics and other training details in  Appendix A.
@@ 4.4 Generalized to Untrained Tasks
4.5 Discussion
More than just for instruction fine-tuning, we believe abstract image comprehension capabilities can be enhanced through various strategies:
Designing More Versatile Visual Encoders: First, we observe that current LMMs have weak visual representation of abstract images, which may be caused by visual encoders. Most of them use clip-based encoders, which emphasize semantic features while neglecting purely visual features. It’s promising to explore another visual encoder to enhance understanding of abstract images.
Increasing Image Resolution: Then we observed that most LMMs resize the original image to a resolution of 336x336, as it reduces the number of visual tokens. However, for these abstract images composed of lines and geometric shapes, lowering the resolution results in the loss of a significant amount of geometric features. Increasing image resolution may be a good solution.
@@ 5 Conclusions
Our multi-modal self-instruct can synthesize a vast amount of abstract images and reasoning instructions for LLMs. However, we want to highlight that there remain some limitations or areas for improvement: 1. Our data synthesis process relies on the code generation and reasoning capabilities of LLMs, which are only available in some strong LLMs like GPT-4. Using these models is costly. It is promising to employ some advanced open-source LLMs, e.g., Qwen2, LLama3.2, to synthesize data. 2. This work used code to synthesize abstract images in eight scenarios. In the future, we can expand to more scenarios, thereby producing a massive amount of data.
6 Acknowledgments
We appreciate the support of Kerui Zhang for the human evaluation of our synthetic data and website design of our project.
This work is supported by the "Pioneer" and "Leading Goose" R&D Programs of Zhejiang (No. 2024C03255), the National Natural Science Foundation of China (No. 62376245), the Fundamental Research Funds for the Central Universities (226-2024-00170), the project of the Donghai Laboratory (Grant no. DH-2022ZY0013), National Key Research and Development Project of China (No. 2018AAA0101900), and MOE Engineering Research Center of Digital Library.
@@ D.3 Data Synthesis
Data synthesis is widely used in LLM training to supplement the insufficiency of instruction-following data. Many studies focus on generating high-quality synthetic data either distilling dialogue data from a strong LLM (Wang et al., 2022b; Zhang et al., 2022, 2023c; Xu et al., 2023a; Yu et al., 2023a; Chen et al., 2023a; Zhao et al., 2023), or using external tools to refine LLM-generated synthetic data (Wei et al., 2023; Lee et al., 2024). For instance, Wang et al. (2022b) proposed Self-Instruct to improve the instruction-following ability of LLMs via their own generation of instruction data. Xu et al. (2023a) further generated more complex instruction through Evol-Instruct. Yu et al. (2023a) synthesized a mathematical dataset from LLMs by bootstrapping mathematical questions and rewriting the question from multiple perspectives. Wei et al. (2023) can generate diverse and realistic coding problems from open-source code snippets. Lei et al. (2024) can also create high-quality large code datasets for LLMs. It simulates programmers writing code and conducting unit tests through agent interactions, ensuring annotation accuracy with an external code executor.